OpenStudy (lilmike234):

Do the point (3,0) lie on the graph of P(x)=x^3-7x^2+15x-9 ?

Vocaloid (vocaloid):

let x = 3 and evaluate P(x)=x^3-7x^2+15x-9

OpenStudy (lilmike234):

I came up with -18 so it doesn't lie on this graph ?

Vocaloid (vocaloid):

hm, check your math one more time

Vocaloid (vocaloid):

3^3 = 27 7(3)^2 = 63 15(3) = 45 27-63+45-9 = ?

OpenStudy (lilmike234):

ooo so I would get 0 ?

Vocaloid (vocaloid):

yes

OpenStudy (lilmike234):

So (3,0) doesn't lie on the graph ?

Vocaloid (vocaloid):

think carefully (3,0) means that P = 0 when x = 3, so (3,0) lies on the graph